OpenStudy (anonymous):

Please help. i have a test today and I'm not understand some of these problems. The line passes through (-6,1) and is perpendicular to the line whose equation is x+3y=9 (thank you:))

OpenStudy (anonymous):

i got y=1/3x+3 but the answer is y=3x+19

OpenStudy (anonymous):

do you know what the slope of the line \[x+3y=9\] is? that is the first step to finding the answer

OpenStudy (anonymous):

oh I'm sorry, i forgot to write that...it's -2

OpenStudy (anonymous):

actually it is \[-\frac{1}{3}\]

OpenStudy (anonymous):

\[ax+by=c\] slope is \[-\frac{a}{b}\]

OpenStudy (anonymous):

im sorry...yes, youre right..i was looking at a different problem (so out of it)

OpenStudy (anonymous):

so since the slope is \[-\frac{1}{3}\] perpendicular line will have slope 3

OpenStudy (anonymous):

that should get you started for the right answer

OpenStudy (anonymous):

and now i use point slope form with the -1/3 as the slope part?

OpenStudy (anonymous):

wait, dont i have to change the x+3y=9 to slope intercept form?